A machine costs $8,000 and will have a useful life of 4 years and has a scrap
value of $3200.
Assuming straight line depreciation, write an equation to give the book value
of the machine after n number of years

Answer by Theo(13342)

the formula for straight line depreciation is:

(cost minus scrap value) / number of years.

Your depreciation calculation is therefore:

annual depreciation = (8,000 - 3,200) / 4 = 4,800 / 4 = 1,200 per year.

at the end of n years, the book value will be original cost minus depreciation * number of years.

this can be shown as b = o - d*n
b = book value
o = original value
d = depreciation per year
n = number ofyears

at the end of year 0 = 8,000 ***** = original price of vehicle.
at the end of year 1 = 6,800
at the end of year 2 = 5,600
at the end of year 3 = 4,400
at the end of year 4 = 3,200 ***** = scrap value of vehicle.